General Journal entry form:

June 1 . Purchased equipment in the amount of $ 40,000 by signing a one-year note with a stated interest rate of 6% per year. The interest and principal of the note will be paid off on the maturity date (5/31/2021).
2 . Prepaid $ 600 cash for four-month insurance (from 6/1/2020 to 9/30/2020).

8. Provided plumbing services of $5,600 on account.

13 . Received $ 4,300 cash in advance for maintenance services to be performed in the future.
17. Purchased $ 1,780 of supplies on account.
22. Collected $ 1,200 cash from accounts receivable.

30. Accrued unpaid June salaries of $ 3,600
30. Counted supplies and determined that $ 420 of supplies (purchased on June 17) remain on hand at the end of June. Adjusted for supplies expense.

30. Adjusted for June insurance expense (see June 2).
30. Accrued June interest on the note (see June 1).
Answer #1

Answer:

Credit Date June. 1 Debit $40,000 General Journal Equipment Notes Payable (to record purchase of equipment signing note) $40,

Calculations:

Supplies expense = 1780 - 420 = $1,360

Insurance expense for June = 600 / 4 = $150

Interest expense for June = 40000 * 6%/12 = $200